40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 6 17:15:15 2012, Jenne, Update, ASS, Filter banks not working 
    Reply  Tue Aug 7 00:22:10 2012, Jenne, Update, ASS, Filter banks working 
       Reply  Thu Aug 9 12:31:36 2012, Jenne, Update, ASS, ASS rebuilt again 
Message ID: 7135     Entry time: Thu Aug 9 12:31:36 2012     In reply to: 7099
Author: Jenne 
Type: Update 
Category: ASS 
Subject: ASS rebuilt again 
I was (in between Eric's measurements) driving the YARM ASS dithers, and noticed that instead of driving ITMY PIT, I was driving ITMX PIT. I looked in the model, and when I re-did the model after an svn revert a few days ago, it looks like I got the shmem parts for the ITM PIT signals backwards. I have fixed that, rebuilt, installed and restarted the ass model.
ELOG V3.1.3-